OpenStudy (anonymous):

Jenna has a bag that contains 7 red marbles and 25 blue marbles. She selects a marble at random, and then, without replacing the first one, selects another marble at random. What is the probability that Jenna selects a blue marble and then a red marble? Round your answer to the nearest percent. P(blue and red) ≈ %

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Didn't we already do this one? You're still not showing your work. Total Marbles = 32 p(blue marble on first draw) = What / 32?

OpenStudy (anonymous):

0.78?

OpenStudy (anonymous):

@tkhunny is 0.78 right?

OpenStudy (anonymous):

@hysusonic

OpenStudy (tkhunny):

Well, I was hoping for 25/32. It's more instructive than 0.78125 Okay, we pull out one blue ball. What is p(red ball on second draw) = What/31?

OpenStudy (anonymous):

7/32=0.21

OpenStudy (tkhunny):

Nope. We took out a blue ball. Try again.

OpenStudy (anonymous):

7/31?

OpenStudy (tkhunny):

Perfect. So, what is the probability of BOTH of those events? p(blue then red) = p(blue on 1st)*p(red on second) = What?

OpenStudy (anonymous):

i don't understand

OpenStudy (tkhunny):

It's just substitution problem at this point. p(blue then red) = p(blue on 1st)*p(red on second) = \(\dfrac{25}{32}\cdot\dfrac{7}{31}\) That's all there is to it.

OpenStudy (anonymous):

i got 175/992

OpenStudy (anonymous):

17.6?

OpenStudy (tkhunny):

There it is. The fractions are just for learning what is going on. 17.6% or 0.176